### v1

@@ 3. $a=3$
Let $u_{long}=0^{x}1^{\alpha_{1}}0^{x}1^{\alpha_{2}}$ and $v_{long}=0^{x}1^{\beta_{1}}0^{x}1^{\beta_{2}}$, where $\alpha_{1}+\alpha_{2}=\beta_{1}+\beta_{2}=n_{1}$. Assume without loss of generality that $\alpha_{1}\leqslant\beta_{1}\leqslant\beta_{2}\leqslant\alpha_{2}$. Let
[…]
@@
where $r_{1}+\ldots+r_{i}=\alpha_{2}$, $r_{i+1}+\ldots+r_{l}=\alpha_{1}$, $p_{1}+\ldots+p_{j}=\beta_{2}$, $p_{j+1}+\ldots+p_{l}=\beta_{1}$. Recall that $y_{u}$ and $y_{v}$ are the lengths of the second longest blocks of 00’s in $u$ and $v$, i.e. the largest length smaller than $x$. Let $y_{max}=\max(y_{u},y_{v})$. Consider two cases.
(a) $\beta_{1}>\alpha_{1}$
Remark 4.6.
The case when both words are special with two blocks $0^{x}$ and $dist(u_{long})\neq dist(v_{long})$ satisfies this case 3(a). We will use it in case 1 of Lemma 4.7.
@@
Since $\beta_{1}>\alpha_{1}$, then $\alpha_{1}<\beta_{1}\leqslant\beta_{2}<\alpha_{2}$. Then the word $u^{\prime}=0^{y_{max}+1}1^{\alpha_{1}}0^{y_{max}+1}1^{\alpha_{2}}$ is a subword of the word $u$ and is not a subword of $v$, and $|u^{\prime}|\leqslant n_{1}+2y_{max}+2$. Then either we have a required subword, or $n_{1}+2y_{max}>\frac{3}{4}n+2$. Now we assume that $n_{1}+2y_{max}>\frac{3}{4}n+2$.
If $i\neq j$, then the word $u^{\prime\prime}=0^{y_{max}+1}(01)^{i}0^{y_{max}+1}(01)^{l-i}$ is a subword of the word $u$ and is not a subword of $v$ and $|u^{\prime\prime}|\leqslant 2l+y_{max}$. If $42l+y_{max}>\frac{3}{4}n+$ 4, then
$(2l+2y_{max})+2(n_{1}+2y_{max})>\frac{3}{4}n+4+2(\frac{3}{4}n+2),$
which implies $82n_{1}+2l+6y_{max}>\frac{9}{4}n+$ 8. Since $n_{0}>2x+y_{max}+(l-3)$, then
$2n=2n_{1}+2n_{0}>2n_{1}+2(2x+y_{max}+(l-3))>2n_{1}+2l+6y_{max}-6>\frac{9}{4}n+2,$
@@
a contradiction. So, $|u^{\prime}|\leqslant\frac{3}{4}n+4$, and hence in the case $i\neq j$, we have a required subword. Now assume that $i=j$.
@@
Since $\alpha_{1}<\beta_{1}$ and $\alpha_{2}>\beta_{2}$, then there are indices $t_{1}\leqslant i$ and $t_{2}\leqslant l-i$ such that $r_{t_{1}}<p_{t_{1}}$ and $r_{i+t_{2}}>p_{i+t_{2}}$. Assume without loss of generality that $r_{t_{1}}<r_{i+t_{2}}$. Consider the subword $v^{\prime}=0^{y_{max}+1}(10)^{t_{1}-1}1^{r_{t_{1}}+1}(01)^{l-t_{1}}$ of $v$. Then there is at most one overlay by 1’s of $v$ on $u$ such that $v^{\prime}$ coincides with an equal word (modulo selection of 0’s and 1’s from each block). Since $u\neq v$, then for this overlay by 1’s either there is a block $0^{s_{1}}$ in $u$ which is smaller than the corresponding block in $u$ ($s_{1}\leqslant y_{max}$) or there is a block $1^{s_{2}}$ in $v$ which is bigger than the corresponding block in $u$. In the first case we add to $v^{\prime}$ the block $0^{s_{1}+1}$ to the corresponding place; we let $v_{1}$ denote the obtained subword of $v$. In the second case we add to $v^{\prime}$ the block $1^{s_{2}}$ to the corresponding place; we let $v_{2}$ denote the obtained subword of $v$. The words $v_{1}$ and $v_{2}$ are subwords of the word $v$ and are not subwords of $u$. We have that $|v_{1}|\leqslant 2l+2y_{max}+r_{t_{1}}$ and $|v_{2}|\leqslant n_{1}+l+y_{max}$.
@@
a contradiction. Then $|v_{2}|\leqslant\frac{3}{4}n+4$.
Since $(r_{t_{1}}-1)+(r_{i+t_{2}}-1)\leqslant n_{1}-l$, then
@@
$\left(\frac{n_{1}}{2}+\frac{3}{2}l+2y_{max}\right)+(n_{1}+2y_{max})<\frac{3}{2}(n_{1}+2x+y_{max}+(l-3))+\frac{9}{2}<\frac{3}{2}n+5,$
a contradiction. Then $|v_{1}|\leqslant\frac{3}{4}n+4$.
Since for the lengths of the obtained subwords in both cases we have $|v_{2}|\leqslant\frac{3}{4}n+4$ and $|v_{1}|\leqslant\frac{3}{4}n+4$, we have a proof of the lemma in case (a).
(b) $\beta_{1}=\alpha_{1}$.
Consider an overlay by 1’s the word $v$ on $u$ such that $u_{long}$ and $v_{long}$ coincide. There is only one such overlay, since $\alpha_{1}<\frac{n_{1}}{2}$ ($u$ is not special). Since $u\neq v$, there are two blocks $0^{s}$ and $0^{t}$ in $u$ such that $0^{s}$ is bigger than the corresponding block in $v$ and $0^{t}$ (possible empty) is smaller than the corresponding block in $v$ ($s,t\leqslant y$). Without loss of generality assume that $s>t$ (in the case $s\leqslant t$ the proof is similar). We now consider the word $v_{1}=1^{n_{1}}$ and add $y_{max}+1$ zeros from each of block $0^{x}$ to $v_{1}$ to the corresponding places. We also add $t+1$ zeros to $v_{1}$ from the block which is bigger than corresponding block $0^{t}$ in $u$. We let $v_{1}^{\prime}$ denote the obtained subword of $v$. The word $v_{1}^{\prime}$ is a subword of the word $v$ and is not a subword of $u$, and $|v_{1}^{\prime}|=n_{1}+2(y_{max}+1)+(t+1)$. If $n_{1}+2y_{max}+t\leqslant\frac{3}{4}n+2$, then we have a required subword, so it remains to considerthe case $n_{1}+2y_{max}+t>\frac{3}{4}n+2$.
Consider the subword $v_{2}=0^{y_{max}+1}1(01)^{l-1}$ of $v$. There are two overlays by 1’s of $v$ on $u$ such that $v_{2}$ coincides with an equal subword (modulo selection of 0’s and 1’s from each block). For each overlay there exists either a block of 00’s or a block of $1$’s in $u$ which is smaller than the corresponding block in $v$ and vice versa. Consider three cases.
$\bullet$ If for both overlays there are blocks of 1’s in $v$ which are bigger than the corresponding blocks of 1’s in $u$, then we add them to $v_{2}$. We let $v_{2}^{\prime}$ denote the obtained subword word of $v$. The word $v_{2}^{\prime}$ is the subword of the word $v$ and is not a subword of $u$, and $|v_{2}^{\prime}|\leqslant n_{1}+l+y_{max}$, since $v_{2}^{\prime}$ has at most $n_{1}$ occurrences of 1’s and $l+y_{max}$ occurrences of 0’s. If $n_{1}+l+y_{max}>\frac{3}{4}n+4$, then
@@
a contradiction. Then $|v_{2}^{\prime}|<\frac{3}{4}n+4$. In this case the lemma is proved.
$\bullet$ If for both overlays there are blocks $0^{s_{1}}$ and $0^{s_{2}}$ in $u$ which are smaller than the corresponding blocks of 0’s in $v$, then we add blocks $0^{s_{1}+1}$ and $0^{s_{2}+1}$ to $v_{2}$ to the corresponding places. We let $v_{2}^{\prime\prime}$ denote the obtained subword of $v$. The word $v_{2}^{\prime\prime}$ is the subword of the word $v$ and is not a subword of $u$ and $|v_{2}^{\prime\prime}|\leqslant 2l+3y_{max}$ since $|v_{2}|=2l+y_{max}$ and we add $s_{1}+s_{2}\leqslant 2y_{max}$ occurrences of 0’s. If $42l+3y_{max}>\frac{3}{4}n+$ 4, then
@@
a contradiction. Then $|v_{2}^{\prime\prime}|<\frac{3}{4}n+4$. In this case the lemma is proved.
$\bullet$ If for one overlay there is a block of 1’s in $v$ which is bigger than the corresponding block of 1’s in $u$ and for the other overlay there is a block of 0’s in $u$ which is smaller than the corresponding block of 0’s in $v$, then for the first overlay there are blocks $1^{t_{1}}$ and $1^{t_{2}}$ in $u$ such that $1^{t_{1}}$ is smaller than the corresponding block in $v$ and $1^{t_{2}}$ is bigger than the corresponding block in $v$. Assume without loss of generality that $t_{1}<t_{2}$. Let $0^{s}$ be the block of 0’s in $u$ which is smaller than the corresponding block in $v$ for the second overlay. We add $t_{1}$ occurrences of 1’s and $s$ occurrences of 0’s to $v_{2}$. We let $v_{2}^{\prime\prime\prime}$ denote the obtained subword word of $v$. The word $v_{2}^{\prime\prime\prime}$ is a subword of the word $v$ and is not a subword of $u$, and $|v_{2}^{\prime\prime\prime}|\leqslant 2l+\frac{n_{1}-l}{2}+2y_{max}$, since $|v_{2}|=2l+y_{max}$ and we added at most $\frac{n_{1}-l}{2}$ occurrences of 1’s and and at most $y_{max}$ occurrences of 0’s. If $42l+\frac{n_{1}-l}{2}+2y_{max}>\frac{3}{4}n+$ 4, then
@@
a contradiction. Then $|v_{2}^{\prime\prime\prime}|<\frac{3}{4}n+4$. In this case the lemma is proved.
4. $(a,b)=(2,\geqslant 3)$.
Let $u_{long}=0^{x}1^{\alpha_{1}}0^{x}1^{\alpha_{2}}$ (with $\alpha_{1}\leqslant\alpha_{2}$) and $v_{long}=0^{x}1^{\beta_{1}}0^{x}1^{\beta_{2}}\ldots 0^{x}1^{\beta_{b}}$. Since $u_{long}$ is not special and $b\geqslant 3$, then $\alpha_{2}>\frac{n}{2}>\alpha_{1}$ and $\beta_{i}\leqslant\frac{n}{2}$ for each $1\leqslant i\leqslant b$. If there is $\beta_{i}\neq\alpha_{1}$ then we can prove lemma as in 3(a) for the words $u_{long}=0^{x}1^{\alpha_{1}}0^{x}1^{\alpha_{2}}$ and $v_{long}=0^{x}1^{\beta_{i}}0^{x}1^{n_{1}-\beta_{i}}$, since $(\alpha_{1},\alpha_{2})\neq(\beta_{i},n_{1}-\beta_{i})$. Assume that $\beta_{i}=\alpha_{1}$ for each $i$. Since $u$ is not special, then $\alpha_{1}\neq\frac{n_{1}}{3}$. Then $b\geqslant 4$. So, we can prove the lemma as in the Case 3(a) for the words $u_{long}=0^{x}1^{\alpha_{1}}0^{x}1^{\alpha_{2}}$ and $v_{long}=0^{x}1^{2\beta_{1}}0^{x}1^{n_{1}-2\beta_{1}}$ since $(\alpha_{1},\alpha_{2})\neq(2\beta_{1},n_{1}-2\beta_{1})$.
The lemma is proved. ∎
@@
Let $u_{long}$ and $v_{long}$ be special words. Then there is a subword of only one of the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
Proof. Recall that a special word has at least two blocks of 0’s, that is, $a\geqslant 2$ and $b\geqslant 2$. Recall that $dist(u_{long})$ denotes the length of small blocks of 1’s in a special word.
First we prove this lemma in the case $dist(u_{long})\neq dist(v_{long})$:
Case 1. $dist(u_{long})\neq dist(v_{long})$.
The case $a=b=2$ has been proved in Lemma 4.5. Now assume that $a\geqslant 3$.
$\bullet$ $a\geqslant 3$ and $b\geqslant 3$.
Assume that $dist(u_{long})>dist(v_{long})$. In this case we can take the following word $v_{1}$, which is a subword of $v$ and not a subword of $u$: the word $v_{1}$ contains $n_{1}$ occurrences of 1 and two blocks $0^{y+1}$ at a distance of $dist(v_{long})$. Since $a\geqslant 3$, then
@@
$\bullet$ $a\geqslant 3$ and $b=2$ (or, symmetrically, $a=2$ and $b\geqslant 3$).
Without loss of generality consider the case $a\geqslant 3$ and $b=2$. In this case $dist(u_{long})<dist(v_{long})$, since $dist(u_{long})\leqslant\frac{n_{1}}{3}$ and $dist(v_{long})\geqslant\frac{n_{1}}{3}$ (the latter inequality comes from the fact that for $b=2$ we have either $v_{long}=0^{x}1^{\frac{n_{1}}{2}}0^{x}1^{\frac{n_{1}}{2}}$ or $v_{long}=0^{x}1^{\frac{n_{1}}{3}}0^{x}1^{\frac{2n_{1}}{3}}$, which gives in the first case $dist(v_{long})=\frac{n_{1}}{2}$ and in the second case $dist(v_{long})=\frac{n_{1}}{3}$). Consider two subwords of $u$: a subword $u_{1}$ containing $n_{1}$ occurrences of 1 and two blocks $0^{x}$ at the distance $dist(u_{long})$ and a subword $u_{2}$ containing one zero and one one from every block and three blocks $0^{x}$. Both words $u_{1}$ and $u_{2}$ are subwords of the word $u$ and is not a subword of $v$. So either lemma proved or $|u_{1}|=n_{1}+2x\geqslant\frac{3}{4}n+5$ and $|u_{2}|=2l+3(x-1)\geqslant\frac{3}{4}n+5$. Summing the second inequality with the tripled first inequality we get:
$203(n_{1}+2x)+(2l+3x-3)\geqslant 3n+$ 20 .
@@
From now on we will assume that $dist(u_{long})=dist(v_{long})=k$. Then $u_{long}$ and $v_{long}$ are of the form
$0^{x}1^{k}\ldots 0^{x}1^{k}(0^{x}|\varepsilon)1^{k}0^{x}1^{k}\ldots 0^{x}1^{k}(0^{x}|\varepsilon)1^{k}0^{x}1^{k}\ldots 0^{x}1^{k},$
@@
where the notation $(0^{x}|\varepsilon)$ is a standard notation from regular expression meaning that we either take $0^{x}$ or the empty word $\varepsilon$. So, if $u_{long}$ or $v_{long}$ is of type 1, then we choose $0^{x}$ in both places; if $u_{long}$ or $v_{long}$ is of type 2, then we choose $0^{x}$ in one of the places and $\varepsilon$ in the other place; if $u_{long}$ or $v_{long}$ is of type 3, then we choose $\varepsilon$ in both places. We now add to $u_{long}$ and $v_{long}$ blocks of 0’s from $u$ and $v$ which are at positions of $(0^{x}|\varepsilon)$. These blocks are either $0^{x}$ (if we had $0^{x}$ in $u_{long}$ or symmetrically $v_{long}$), or they can be shorter (if we had $\varepsilon$ in $u_{long}$ or symmetrically $v_{long}$). The new subwords look like
$0^{x}1^{k}\ldots 0^{x}1^{k}0^{\alpha_{u}}1^{k}0^{x}1^{k}\ldots 0^{x}1^{k}0^{\beta_{u}}1^{k}0^{x}1^{k}\ldots 0^{x}1^{k},$
$0^{x}1^{k}\ldots 0^{x}1^{k}0^{\alpha_{v}}1^{k}0^{x}1^{k}\ldots 0^{x}1^{k}0^{\beta_{v}}1^{k}0^{x}1^{k}\ldots 0^{x}1^{k},$

### v2

@@ 4.2 Lemmas constituting the proof
Let $u_{\operatorname{long}}=0^{x}1^{\alpha_{1}}0^{x}1^{\alpha_{2}}$ and $v_{\operatorname{long}}=0^{x}1^{\beta_{1}}0^{x}1^{\beta_{2}}$, where $\alpha_{1}+\alpha_{2}=\beta_{1}+\beta_{2}=n_{1}$. Assume without loss of generality that $\alpha_{1}\leqslant\beta_{1}\leqslant\beta_{2}\leqslant\alpha_{2}$. Let
$u=0^{x}\,\,1^{r_{1}}0^{\gamma_{1}}\cdots 0^{\gamma_{i-1}}1^{r_{i}}\,\,0^{x}\,\,1^{r_{i+1}}0^{\gamma_{i}}\cdots 0^{\gamma_{l-2}}1^{r_{l}}$
$v=0^{x}\,\,1^{p_{1}}0^{\phi_{1}}\cdots 0^{\phi_{j-1}}1^{p_{j}}\,\,0^{x}\,\,1^{p_{j+1}}0^{\phi_{j}}\cdots 0^{\phi_{l-2}}1^{p_{l}},$
@@
where $r_{1}+\ldots+r_{i}=\alpha_{2}$, $r_{i+1}+\ldots+r_{l}=\alpha_{1}$, $p_{1}+\ldots+p_{j}=\beta_{2}$, $p_{j+1}+\ldots+p_{l}=\beta_{1}$. Recall that $y_{u}$ and $y_{v}$ are the lengths of the second longest blocks of 0’s in $u$ and $v$, i.e. the largest length smaller than $x$. Let $y_{max}=\max(y_{u},y_{v})$. Consider two cases.
Case 3.1. $\beta_{1}>\alpha_{1}$.
@@
Since $\beta_{1}>\alpha_{1}$, we have $\alpha_{1}<\beta_{1}\leqslant\beta_{2}<\alpha_{2}$. Then the word $u^{\prime}=0^{y_{max}+1}1^{\alpha_{1}}0^{y_{max}+1}1^{\alpha_{2}}$ is a subword of the word $u$ and is not a subword of $v$, and $|u^{\prime}|\leqslant n_{1}+2y_{max}+2$. Then either we have a required subword, or $n_{1}+2y_{max}>\frac{3}{4}n+2$. Now we assume that $n_{1}+2y_{max}>\frac{3}{4}n+2$.
If $i\neq j$, then the word $u^{\prime\prime}=0^{y_{max}+1}1(01)^{i-1}0^{y_{max}+1}(01)^{l-i}$ is a subword of the word $u$ and is not a subword of $v$ and $|u^{\prime\prime}|\leqslant 2l+y_{max}$. If $2l+y_{max}>\frac{3}{4}n+4$, then
$(2l+2y_{max})+2(n_{1}+2y_{max})>\frac{3}{4}n+4+2\left(\frac{3}{4}n+2\right),$
which implies $2n_{1}+2l+6y_{max}>\frac{9}{4}n+8$. Since $n_{0}>2x+y_{max}+(l-3)$, we have
$2n=2n_{1}+2n_{0}>2n_{1}+2(2x+y_{max}+(l-3))>2n_{1}+2l+6y_{max}-6>\frac{9}{4}n+2;$
@@
a contradiction. So, $|u^{\prime}|\leqslant\frac{3}{4}n+4$, and hence in the case $i\neq j$ we have a required subword. Now assume that $i=j$.
@@
Since $\alpha_{1}<\beta_{1}$ and $\alpha_{2}>\beta_{2}$, there are indices $t_{1}\leqslant i$ and $t_{2}\leqslant l-i$ such that $r_{t_{1}}<p_{t_{1}}$ and $r_{i+t_{2}}>p_{i+t_{2}}$. Assume without loss of generality that $r_{t_{1}}<r_{i+t_{2}}$. Consider the subword $v^{\prime}=0^{y_{max}+1}(10)^{t_{1}-1}1^{r_{t_{1}}+1}(01)^{l-t_{1}}$ of $v$. Then there is at most one 1-overlay of $v$ on $u$ such that $v^{\prime}$ coincides with an occurrence of an equal word (modulo selection of 0’s and 1’s from each block). Since $u\neq v$, for this 1-overlay either there is a block $0^{s_{1}}$ in $u$ which is smaller than the corresponding block in $v$ ($s_{1}\leqslant y_{max}$) or there is a block $1^{s_{2}}$ in $u$ which is bigger than the corresponding block in $v$. In the first case we add to $v^{\prime}$ the block $0^{s_{1}+1}$ to the corresponding place; we let $v_{1}$ denote the obtained subword of $v$. In the second case we add to $v^{\prime}$ the block $1^{s_{2}}$ to the corresponding place; we let $u_{1}$ denote the obtained subword of $u$. The words $v_{1}$ and $u_{1}$ are distinguishing subwords for $u$ and $v$. We have that $|v_{1}|\leqslant 2l+2y_{max}+r_{t_{1}}$ and $|u_{1}|\leqslant n_{1}+l+y_{max}$.
@@
a contradiction. Thus $|u_{1}|\leqslant\frac{3}{4}n+4$.
Since $(r_{t_{1}}-1)+(r_{i+t_{2}}-1)\leqslant n_{1}-l$, we have
@@
$\left(\frac{n_{1}}{2}+\frac{3}{2}l+2y_{max}\right)+(n_{1}+2y_{max})<\frac{3}{2}(n_{1}+2x+y_{max}+(l-3))+\frac{9}{2}<\frac{3}{2}n+5;$
a contradiction. Thus $|v_{1}|\leqslant\frac{3}{4}n+4$.
Since for the lengths of the obtained subwords in both cases we have $|u_{1}|\leqslant\frac{3}{4}n+4$ and $|v_{1}|\leqslant\frac{3}{4}n+4$, we have a proof of the lemma in this case.
Case 3.2. $\beta_{1}=\alpha_{1}$.
Consider a 1-overlay of $v$ on $u$ such that $u_{\operatorname{long}}$ and $v_{\operatorname{long}}$ coincide. There is only one such overlay, since $\alpha_{1}<\frac{n_{1}}{2}$ ($u_{\operatorname{long}}$ is not special). Since $u\neq v$, there are two blocks $0^{s}$ and $0^{t}$ in $u$ such that $0^{s}$ is bigger than the corresponding block in $v$ and $0^{t}$ (possible empty) is smaller than the corresponding block in $v$ ($s,t\leqslant y$). Without loss of generality assume that $s>t$ (in the case $s\leqslant t$ the proof is similar). We now consider the word $v_{1}=1^{n_{1}}$ and add $y_{max}+1$ zeros from each of the block $0^{x}$ to $v_{1}$ to the corresponding places. We also add $t+1$ zeros to $v_{1}$ from the block which is bigger than the corresponding block $0^{t}$ in $u$. We let $v_{1}^{\prime}$ denote the obtained subword of $v$. The word $v_{1}^{\prime}$ is a subword of the word $v$ and is not a subword of $u$, and $|v_{1}^{\prime}|=n_{1}+2(y_{max}+1)+(t+1)$. If $n_{1}+2y_{max}+t\leqslant\frac{3}{4}n+2$, then we have a required subword, so it remains to consider the case $n_{1}+2y_{max}+t>\frac{3}{4}n+2$.
Consider the subword $v_{2}=0^{y_{max}+1}1(01)^{l-1}$ of $v$. There are two 1-overlays of $v$ on $u$ such that $v_{2}$ coincides with an occurrence of an equal subword (modulo selection of 0’s and 1’s from each block). For each 1-overlay there exists either a block of 0’s or a block of 1’s in $u$ which is smaller than the corresponding block in $v$ and vice versa. Consider three cases.
Case 3.2.1. If for both 1-overlays there are blocks of 1’s in $v$ which are bigger than the corresponding blocks of 1’s in $u$, then we add them to $v_{2}$. We let $v_{2}^{\prime}$ denote the obtained subword of $v$. The word $v_{2}^{\prime}$ is a subword of the word $v$ and is not a subword of $u$, and $|v_{2}^{\prime}|\leqslant n_{1}+l+y_{max}$, since $v_{2}^{\prime}$ has at most $n_{1}$ occurrences of 1 and $l+y_{max}$ occurrences of 0. If $n_{1}+l+y_{max}>\frac{3}{4}n+4$, then
@@
a contradiction. Thus $|v_{2}^{\prime}|<\frac{3}{4}n+4$. In this case the lemma is proved.
Case 3.2.2. If for both 1-overlays there are blocks $0^{s_{1}}$ and $0^{s_{2}}$ in $u$ which are smaller than the corresponding blocks of 0’s in $v$, then we add blocks $0^{s_{1}+1}$ and $0^{s_{2}+1}$ to $v_{2}$ to the corresponding places. We let $v_{2}^{\prime\prime}$ denote the obtained subword of $v$. The word $v_{2}^{\prime\prime}$ is a subword of the word $v$ and is not a subword of $u$ and $|v_{2}^{\prime\prime}|\leqslant 2l+3y_{max}$ since $|v_{2}|=2l+y_{max}$ and we add $s_{1}+s_{2}\leqslant 2y_{max}$ occurrences of 0. If $2l+3y_{max}>\frac{3}{4}n+4$, then
@@
a contradiction. Thus $|v_{2}^{\prime\prime}|<\frac{3}{4}n+4$. In this case the lemma is proved.
Case 3.2.3. If for one 1-overlay there is a block of 1’s in $v$ which is bigger than the corresponding block of 1’s in $u$ and for the other 1-overlay there is a block of 0’s in $u$ which is smaller than the corresponding block of 0’s in $v$, then for the first 1-overlay there are blocks $1^{t_{1}}$ and $1^{t_{2}}$ in $u$ such that $1^{t_{1}}$ is smaller than the corresponding block in $v$ and $1^{t_{2}}$ is bigger than the corresponding block in $v$. Assume without loss of generality that $t_{1}<t_{2}$. Let $0^{s}$ be the block of 0’s in $u$ which is smaller than the corresponding block in $v$ for the second 1-overlay. We add $t_{1}$ occurrences of 1 and $s$ occurrences of 0 to $v_{2}$. We let $v_{2}^{\prime\prime\prime}$ denote the obtained subword word of $v$. The word $v_{2}^{\prime\prime\prime}$ is a subword of the word $v$ and is not a subword of $u$, and $|v_{2}^{\prime\prime\prime}|\leqslant 2l+\frac{n_{1}-l}{2}+2y_{max}$, since $|v_{2}|=2l+y_{max}$ and we added at most $\frac{n_{1}-l}{2}$ occurrences of 1 and and at most $y_{max}$ occurrences of 0. If $2l+\frac{n_{1}-l}{2}+2y_{max}>\frac{3}{4}n+4$, then
@@
a contradiction. Thus $|v_{2}^{\prime\prime\prime}|<\frac{3}{4}n+4$. In this case the lemma is proved.
Case 4. $(a,b)=(2,\geqslant 3)$.
In this case we have $u_{\operatorname{long}}=0^{x}1^{\alpha_{1}}0^{x}1^{\alpha_{2}}$ and $v_{\operatorname{long}}=0^{x}1^{\beta_{1}}0^{x}1^{\beta_{2}}\cdots 0^{x}1^{\beta_{b}}$, for some integers $x$, $\beta_{1}$, …, $\beta_{b}$ and $\alpha_{1}\leqslant\alpha_{2}$. Since $u_{\operatorname{long}}$ is not special and $b\geqslant 3$, we have $\alpha_{2}>\frac{n}{2}>\alpha_{1}$ and $\beta_{i}\leqslant\frac{n}{2}$ for each $1\leqslant i\leqslant b$. If there is $\beta_{i}\neq\alpha_{1}$ then the proof is similar to the proof in Case 3.1 for the words $u_{\operatorname{long}}=0^{x}1^{\alpha_{1}}0^{x}1^{\alpha_{2}}$ and $v_{\operatorname{long}}=0^{x}1^{\beta_{i}}0^{x}1^{n_{1}-\beta_{i}}$, since $(\alpha_{1},\alpha_{2})\neq(\beta_{i},n_{1}-\beta_{i})$. Assume that $\beta_{i}=\alpha_{1}$ for each $i$. Since $u$ is not special, we have $\alpha_{1}\neq\frac{n_{1}}{3}$. Then $b\geqslant 4$. So, we can prove the lemma as in the Case 3.1 for the words $u^{\prime}=u_{\operatorname{long}}=0^{x}1^{\alpha_{1}}0^{x}1^{\alpha_{2}}$ and $v^{\prime}=0^{x}1^{2\beta_{1}}0^{x}1^{n_{1}-2\beta_{1}}$ since $(\alpha_{1},\alpha_{2})\neq(2\beta_{1},n_{1}-2\beta_{1})$. ∎
@@
Let $u_{\operatorname{long}}$ and $v_{\operatorname{long}}$ be special words. Then there is a distinguishing subword of length at most $\frac{3}{4}n+4$ for $u$ and $v$.
Proof. We recall that a special word has at least two blocks of 0’s, that is, $a\geqslant 2$ and $b\geqslant 2$, and that $dist(u_{\operatorname{long}})$ denotes the length of the shortest block of 1’s in a special word.
First we prove this lemma in the case $dist(u_{\operatorname{long}})\neq dist(v_{\operatorname{long}})$:
Case 1. $dist(u_{\operatorname{long}})\neq dist(v_{\operatorname{long}})$.
The proof in the case $a=b=2$ is similar to the proof in Case 3.1 of Lemma 4.5. Now assume that $a\geqslant 3$.
Case 1.1. $a\geqslant 3$ and $b\geqslant 3$.
Assume that $dist(u_{\operatorname{long}})>dist(v_{\operatorname{long}})$. In this case we can take the following word $v_{1}$, which is a subword of $v$ and not a subword of $u$: the word $v_{1}$ contains $n_{1}$ occurrences of 1 and two blocks $0^{y+1}$ at distance $dist(v_{\operatorname{long}})$. Since $a\geqslant 3$, we have
@@
Case 1.2. $a\geqslant 3$ and $b=2$ (or, symmetrically, $a=2$ and $b\geqslant 3$).
Without loss of generality, consider the case $a\geqslant 3$ and $b=2$. In this case $dist(u_{\operatorname{long}})<dist(v_{\operatorname{long}})$, since $dist(u_{\operatorname{long}})\leqslant\frac{n_{1}}{3}$ and $dist(v_{\operatorname{long}})\geqslant\frac{n_{1}}{3}$ (the latter inequality comes from the fact that for $b=2$, we have either $v_{\operatorname{long}}=0^{x}1^{\frac{n_{1}}{2}}0^{x}1^{\frac{n_{1}}{2}}$ or $v_{\operatorname{long}}=0^{x}1^{\frac{n_{1}}{3}}0^{x}1^{\frac{2n_{1}}{3}}$, which gives in the first case $dist(v_{\operatorname{long}})=\frac{n_{1}}{2}$ and in the second case $dist(v_{\operatorname{long}})=\frac{n_{1}}{3}$). Consider two subwords of $u$: a subword $u_{1}$ containing $n_{1}$ occurrences of 1 and two blocks $0^{x}$ at distance $dist(u_{\operatorname{long}})$ and a subword $u_{2}$ containing one 0 and one 1 from each block and three blocks $0^{x}$. Both words $u_{1}$ and $u_{2}$ are subwords of the word $u$ and are not subwords of $v$. So either we have a distinguishing subword, or $|u_{1}|=n_{1}+2x\geqslant\frac{3}{4}n+5$ and $|u_{2}|=2l+3(x-1)\geqslant\frac{3}{4}n+5$. Summing the second inequality with the tripled first inequality we get:
$3(n_{1}+2x)+(2l+3x-3)\geqslant 3n+20.$
@@
From now on we assume that $dist(u_{\operatorname{long}})=dist(v_{\operatorname{long}})=k$. Then $u_{\operatorname{long}}$ and $v_{\operatorname{long}}$ are of the form
$0^{x}1^{k}\cdots 0^{x}1^{k}(0^{x}|\varepsilon)1^{k}0^{x}1^{k}\cdots 0^{x}1^{k}(0^{x}|\varepsilon)1^{k}0^{x}1^{k}\cdots 0^{x}1^{k},$
@@
where the notation $(0^{x}|\varepsilon)$ is the standard notation for regular expressions meaning that we either take $0^{x}$ or the empty word $\varepsilon$. So, if $u_{\operatorname{long}}$ or $v_{\operatorname{long}}$ is of type 1, then we choose $0^{x}$ in both places; if $u_{\operatorname{long}}$ or $v_{\operatorname{long}}$ is of type 2, then we choose $0^{x}$ in one of the places and $\varepsilon$ in the other place; if $u_{\operatorname{long}}$ or $v_{\operatorname{long}}$ is of type 3, then we choose $\varepsilon$ in both places. We now add to $u_{\operatorname{long}}$ and $v_{\operatorname{long}}$ blocks of 0’s from $u$ and $v$ which are at positions of $(0^{x}|\varepsilon)$. These blocks are either $0^{x}$ (if we had $0^{x}$ in $u_{\operatorname{long}}$ or analogously $v_{\operatorname{long}}$), or they can be shorter (if we had $\varepsilon$ in $u_{\operatorname{long}}$ or analogously $v_{\operatorname{long}}$). The new subwords are of the following form:
$0^{x}1^{k}\cdots 0^{x}1^{k}0^{\alpha_{u}}1^{k}0^{x}1^{k}\cdots 0^{x}1^{k}0^{\beta_{u}}1^{k}0^{x}1^{k}\cdots 0^{x}1^{k},$
$0^{x}1^{k}\cdots 0^{x}1^{k}0^{\alpha_{v}}1^{k}0^{x}1^{k}\cdots 0^{x}1^{k}0^{\beta_{v}}1^{k}0^{x}1^{k}\cdots 0^{x}1^{k},$